# Mohan Xu1, Kai Li1,2,, Guo Chen1,2 & Xiaolin Hu1,2,3,†

### v1

@@ Abstract
In recent years, much speech separation research has focused primarily on improving model performance. However, for low-latency speech processing systems, high efficiency is equally important. Therefore, we propose a speech separation model with significantly reduced parameters and computational costs: Time-frequency Interleaved Gain Extraction and Reconstruction network (TIGER). TIGER leverages prior knowledge to divide frequency bands and compresses frequency information. We employ a multi-scale selective attention module to extract contextual features, while introducing a full-frequency-frame attention module to capture both temporal and frequency contextual information. Additionally, to more realistically evaluate the performance of speech separation models in complex acoustic environments, we introduce a dataset called EchoSet. This dataset includes noise and more realistic reverberation (e.g., considering object occlusions and material properties), with speech from two speakers overlapping at random proportions. Experimental results showed that models trained on EchoSet had better generalization ability than those trained on other datasets to the data collected in the physical world, which validated the practical value of the EchoSet. On EchoSet and real-world data, TIGER significantly reduces the number of parameters by 94.3% and the MACs by 95.3% while achieving performance surpassing state-of-the-art (SOTA) model TF-GridNet. This is the first speech separation model with fewer than 1 million parameters that achieves performance comparable to the SOTA model.
@@ 1 Introduction
Humans possess the ability to focus on a specific speech signal in noisy environments, a phenomenon known as the “cocktail party effect” (Cherry, 1953). In speech processing, the corresponding challenge is accurately separating different sound sources from mixed audio signals, a task referred to as speech separation. Speech separation is typically used as a preprocessing step for speech recognition, as it helps enhance recognition accuracy (Haykin & Chen, 2005). Consequently, it is crucial to ensure that speech separation not only produces clear and distinct outputs on real-world audios but also meets the demands of low latency and computational efficiency (Divenyi, 2004). Therefore, in practical applications, optimizing the parameters and computational costs of separation models to reduce resource consumption becomes particularly important.
@@
First, when designing a separation model, we should fully take into account the actual application scenarios of the speech processing system, which require low latency and low computational complexity. However, many approaches have primarily focused on improving speech separation performance. For example, TF-GridNet (Wang et al., 2023) utilizes bidirectional LSTMs and self-attention mechanisms in an alternating manner, while SepFormer (Subakan et al., 2021) incorporates Transformer-based architectures to enhance sequential modeling capabilities. These models have achieved good results on benchmark datasets but have large model sizes. To make the model more applicable in computationally constrained real-world scenarios, it is critical to make the model more lightweight while maintaining high performance. TDANet (Li et al., 2023) introduces an efficient lightweight architecture using top-down attention, achieving competitive performance with lower computational costs than SepFormer (Subakan et al., 2021). However, as a time-domain method, TDANet struggles to leverage frequency-domain information. On the other hand, time-frequency domain approaches like TF-GridNet (Wang et al., 2023) model both time and frequency dimensions but require higher computational resources. BSRNN (Luo & Yu, 2023), which is the SOTA model for music separation, reduces the computational burden by focusing on frequency bands, but the band-split strategy is under-explored in speech separation. How to balance computational efficiency and separation quality by effectively utilizing frequency characteristics in speech separation is still a big challenge.
@@
Second, the commonly used speech separation datasets still exhibit a significant gap from real-world scenarios. Many methods relied on the WSJ0-2mix dataset (Hershey et al., 2016) for evaluation, which only contains clean audio without noise or reverberation. In addition, the speech signals from two speakers are fully overlapping. Models trained on this kind of dataset are subject to weak generalization and robustness in real-world environments (Kadıoğlu et al., 2020). Although the WHAMR! dataset (Maciejewski et al., 2020) introduces noise and reverberation to WSJ0-2mix, the generated reverberation fails to fully take into account factors such as object occlusion and material properties, and the diversity of acoustic scenarios remains limited. Therefore, to more accurately train and evaluate speech separation models for practical use, a dataset that more closely resembles real-world environments is necessary. Specifically, this dataset should include different noise types, cover a wide range of realistic acoustic environments, and have speech overlap ratios that are randomly distributed.
To address the aforementioned two issues, we propose a novel time-frequency domain speech separation model called Time-frequency Interleaved Gain Extraction and Reconstruction network (TIGER), along with a speech separation dataset containing reverberation and noise (EchoSet). TIGER employs frequency band-split and alternating time-frequency modeling, resulting in a model with significantly smaller parameters and higher computational efficiency. EchoSet features diverse background noise and realistic reverberation, with the shape and material of objects in the room specially considered in reverberation generation. Besides, we select diverse scenarios and add speaker and microphone constraints to simulate real-life acoustic scenarios.
Our main contributions are as follows:
@@
Experiments show that models trained on EchoSet generalized better on real-world data than those trained on benchmark dataset LRS2-2Mix (Li et al., 2023) and Libri2Mix (Cosentino et al., 2020), validating that audio in EchoSet are closer to the physical world. We then comprehensively evaluated TIGER on Libri2Mix, LRS2-2Mix and EchoSet. As the dataset becomes more complex, TIGER’s superiority in performance becomes more significant. On EchoSet, which is the most complicated among the three datasets, TIGER improved the performance by about 5% compared with TF-GridNet, while reducing the parameters and MACs by 94.3% and 95.3% respectively. When tested on real-world data, TIGER also achieved the best separation performance. The remarkable result shows that TIGER provides a new solution for the design of lightweight speech separation models for practical use in the time-frequency domain.
@@ 2 Related Work
Speech separation. Speech separation methods can be divided into time domain and time-frequency domain. Time domain methods directly process the original audio signal. Conv-TasNet (Luo & Mesgarani, 2019) is the cornerstone of time domain separation models. It extracts features by temporal convolutional network (Lea et al., 2016). To improve the performance on long sequence data, DPRNN (Luo et al., 2020) divides the temporal sequence into small blocks and alternately performs intra-block and inter-block modeling, which becomes a common paradigm for many following works (Wang et al., 2023; Subakan et al., 2021). The time-frequency domain methods apply a Short-Time Fourier Transform (STFT) to transform the waveform into a joint representation of time and frequency. TF-GridNet (Wang et al., 2023) enhances the temporal context information by a full-band self-attention module. Although TF-GridNet achieved SOTA performance, it involves huge computational costs.
Lightweight models. Some models (Wang et al., 2023; Yang et al., 2022; Subakan et al., 2021) with high computational complexity are difficult to be applied to real-time speech processing on edge devices. To reduce computational costs, TDANet (Li et al., 2023) draws on the attention mechanism of human brains and designs a lightweight structure. TDANet achieved the same performance with only 5% MACs of SepFormer (Subakan et al., 2021), inspiring the design of lightweight separation models. In music separation, BSRNN (Luo & Yu, 2023) uses prior knowledge to split band, performing band merging on less important bands to compress the feature while retaining key frequency band information. BSRNN surpasses the previous music separation model in performance. The band-split strategy is also inspiring for exploring lightweight speech separation models.
Datasets for speech separation. WSJ0-2mix (Hershey et al., 2016) is an early and commonly used speech separation dataset. However, there are two main problems with WSJ0-2mix. First, utterances of different speakers fully overlap during the conversation, which is not the case in real world. Second, audio in WSJ0-2mix contain no noise or reverberation. WHAM! (Wichern et al., 2019) added environmental noises to WSJ0-2mix. To take a step further, WHAMR! (Maciejewski et al., 2020) added reverberation to the audio. However, it only considered rooms with rectangular shapes and ignored the influence of materials and other non-rectangular shapes in the room on reverberation. Libri2Mix (Cosentino et al., 2020) was proposed based on the observation (Kadıoğlu et al., 2020) that the test performance of Conv-TasNet trained on WSJ0-2mix dropped sharply on other separation datasets. The utterances in Libri2Mix were mixed with sparse overlap, and noises were added to the mixed audio, but reverberation was not considered in Libri2Mix. LRS2-2Mix (Li et al., 2023) was mixed by video clips acquired through BBC. The audio was recorded in real acoustic scenarios, thus containing much noise and reverberation. However, due to the different recording environments of the clips, such as the shapes and materials of the room and objects, the reverberation obtained when the clips were directly mixed is still unrealistic.
@@ 3 TIGER
Figure 1: The overall pipeline for TIGER. We focus on scenarios with only two speakers.
@@ 3.1 Overall Pipeline
The TIGER system (Figure 1) can be divided into five main components: the encoder, the band-split module, the separator, the band-restoration module, and the decoder. Specifically, we first use STFT as the encoder to convert the mixed audio signal $\boldsymbol{S}\in\mathbb{R}^{1\times L}$ into its time-frequency representation $\boldsymbol{X}\in\mathbb{C}^{F\times T}$, where $F$ and $T$ represent the number of frequency bins and time frames, respectively. Next, we apply a frequency band-split strategy, dividing the frequency bands into $K$ sub-bands of varying widths based on their importance. Each sub-band is transformed into a uniform channel size $N$ using 1D convolutions, and these are then stacked along the frequency dimension to produce the feature representation $\boldsymbol{Z}\in\mathbb{R}^{N\times K\times T}$. Thirdly, $\boldsymbol{Z}$ serves as the input to the separator, which uses FFI blocks with shared parameters to model the acoustic characteristics of each speaker. Subsequently, the band-restoration module restores the sub-bands to the full frequency range using separator output $\boldsymbol{J}\in\mathbb{R}^{N\times K\times T}$, and the mask for each speaker $\boldsymbol{M}_{i}\in\mathbb{C}^{F\times T}$ is applied element-wise product to $\boldsymbol{X}$, producing the separated representation for each speaker $\boldsymbol{H}_{i}\in\mathbb{C}^{F\times T}$. Finally, the inverse STFT is used to generate the clean speech signal $\bar{\boldsymbol{P}}_{i}\in\mathbb{R}^{1\times L}$ for each speaker.
@@ 3.3 Separator
In the separator, the input feature $\boldsymbol{Z}$ passes sequentially through $B$ frequency-frame interleaved (FFI) blocks with shared parameters, as shown in Figure 2. In each FFI block, the frequency path is first used to extract contextual information between different sub-bands, producing $\boldsymbol{Z}_{b,f}\in\mathbb{R}^{N\times K\times T}$. Next, we feed $\boldsymbol{Z}_{b,f}$ into the frame path to further model the contextual information between different time frames, generating $\boldsymbol{Z}_{b+1}\in\mathbb{R}^{N\times K\times T}$.
@@
3.3.1 MSA module
The MSA module enhances important features through a selective attention mechanism and is divided into three stages: encoding, fusion, and decoding, as shown in Figure 3(a). Taking the MSA module in the frequency path as an example, the input to the module is $\boldsymbol{Z}_{b}$.
The encoding stage. This stage aims to capture multi-scale acoustic features. Specifically, we first use multiple 1D convolutional layers (with a stride of 2 and channel of $H$) to progressively downsample the frequency dimension to $\frac{K}{2^{D}}$, resulting in a set of multi-scale acoustic features $\{\boldsymbol{E}_{d}\in\mathbb{R}^{H\times\frac{K}{2^{d}}\times T}|d=[0,D]\}$, where $d$ denotes the $d$-th layer of downsampling. Next, we apply average pooling layers, denoted as $\lambda(\cdot)$, to downsample all $\boldsymbol{E}_{d}$ to the same frequency resolution $\frac{K}{2^{D}}$. Subsequently, the features with different frequency resolutions are fused into global features $\boldsymbol{G}=\sum_{d=0}^{D}\lambda(\boldsymbol{E}_{d}),\boldsymbol{G}\in\mathbb{R}^{H\times\frac{K}{2^{D}}\times T}$ by summing. Finally, a multi-layer convolutional (MLC) network is used to transform $\boldsymbol{G}$ into $\boldsymbol{G}^{\prime}\in\mathbb{R}^{H\times\frac{K}{2^{D}}\times T}$.
@@ 3.3.1 MSA module
where $\odot$ denotes element-wise multiplication. In this way, $\{\boldsymbol{L}_{d}\in\mathbb{R}^{H\times\frac{K}{2^{d}}\times T}~{}|~{}d=[0,D]\}$ contains both local and global information, which helps the model better extract the acoustic feature in the audio mixture.
The decoding stage. In the $d$-th layer, where $d\in[0,D-1]$, the input consists of the decoding result from the previous layer $d+1$ (denoted as $\boldsymbol{D}_{d+1}\in\mathbb{R}^{H\times\frac{K}{2^{d+1}}\times T}$) and the output $\boldsymbol{L}_{d}$ from the fusion stage at the $d$-th layer. $\boldsymbol{D}_{d+1}$ is processed through the SA module to produce $\boldsymbol{D}_{d}$. Specifically, $\boldsymbol{D}_{d+1}$ is transformed using two 1D convolutions to obtain $\alpha\in\mathbb{R}^{H\times\frac{K}{2^{d+1}}\times T}$ and $\beta\in\mathbb{R}^{H\times\frac{K}{2^{d+1}}\times T}$, while $\boldsymbol{L}_{d}$ is transformed through a 1D convolution to produce $\gamma\in\mathbb{R}^{H\times\frac{K}{2^{d}}\times T}$. We then compute:
@@
where $f$ is defined in equation 5. This formulation integrates the decoding result with the output from the fusion stage to generate the next layer of decoded features. In particular, for the layer where $d=D$, $\boldsymbol{D}_{D}=\boldsymbol{L}_{D}\in\mathbb{R}^{H\times\frac{K}{2^{D}}\times T}$. For the layer where $d=0$, we use one 1D convolution to restore the hidden dimension $H$ in $\boldsymbol{D}_{0}\in\mathbb{R}^{H\times K\times T}$ to the feature dimension $N$, obtaining $\bar{\boldsymbol{Z}}_{b}\in\mathbb{R}^{N\times K\times T}$ as the output of the MSA module. In the MSA module of the frequency path, the frequency dimension $K$ is considered the processing dimension. In the frame path, the time dimension $T$ is considered the processing dimension.

### v2

# TIGER: Time-frequency Interleaved Gain Extraction and Reconstruction for Efficient Speech Separation

@@ Abstract
In recent years, much speech separation research has focused primarily on improving model performance. However, for low-latency speech processing systems, high efficiency is equally important. Therefore, we propose a speech separation model with significantly reduced parameters and computational costs: Time-frequency Interleaved Gain Extraction and Reconstruction network (TIGER). TIGER leverages prior knowledge to divide frequency bands and compresses frequency information. We employ a multi-scale selective attention module to extract contextual features, while introducing a full-frequency-frame attention module to capture both temporal and frequency contextual information. Additionally, to more realistically evaluate the performance of speech separation models in complex acoustic environments, we introduce a dataset called EchoSet. This dataset includes noise and more realistic reverberation (e.g., considering object occlusions and material properties), with speech from two speakers overlapping at random proportions. Experimental results showed that models trained on EchoSet had better generalization ability than those trained on other datasets to the data collected in the physical world, which validated the practical value of the EchoSet. On EchoSet and real-world data, TIGER significantly reduces the number of parameters by 94.3% and the MACs by 95.3% while achieving performance surpassing state-of-the-art (SOTA) model TF-GridNet.
@@ 1 Introduction
Humans possess the ability to focus on a specific speech signal in noisy environments, which is a phenomenon known as the “cocktail party effect” (Cherry, 1953). In speech processing, the corresponding challenge is accurately separating different sound sources from mixed audio signals, a task referred to as speech separation. Speech separation is typically used as a preprocessing step for speech recognition, as it helps enhance recognition accuracy (Haykin & Chen, 2005). Consequently, it is crucial to ensure that speech separation not only produces clear and distinct outputs on real-world audio but also meets the demand of high computational efficiency (Divenyi, 2004).
@@
First, when designing a separation model, we should fully take into account the actual application scenarios of the speech processing system, which require low latency and low computational complexity. However, many approaches have primarily focused on improving speech separation performance. For example, TF-GridNet (Wang et al., 2023) utilizes bidirectional LSTMs and self-attention mechanisms in an alternating manner, achieving good results on benchmark datasets but have large model size. To make the separation model more applicable in computationally constrained real-world scenarios, TDANet (Li et al., 2023) introduces an efficient lightweight architecture using top-down attention, achieving competitive performance with lower computational costs than SepFormer (Subakan et al., 2021). However, as a time-domain method, TDANet struggles to leverage frequency-domain information. On the other hand, time-frequency domain approaches like TF-GridNet (Wang et al., 2023) model both time and frequency dimensions but require higher computational resources. BSRNN (Luo & Yu, 2023), which is the SOTA model for music separation, reduces the computational burden by focusing on important frequency bands. The band-split strategy is enlightening but under-explored in speech separation. How to balance computational efficiency and separation quality in speech separation is still a big challenge.
@@
Second, the commonly used speech separation datasets exhibit a significant gap from real-world scenarios. Many methods relied on the WSJ0-2mix dataset (Hershey et al., 2016) for evaluation, which only contains fully-overlapping audio without noise or reverberation. Models trained on this kind of dataset are subject to weak generalization and robustness in real-world environments (Kadıoğlu et al., 2020; Cosentino et al., 2020). Although the WHAMR! dataset (Maciejewski et al., 2020) adds noise and reverberation to WSJ0-2mix, the generated reverberation fails to fully take into account factors such as object occlusion and material properties, and the diversity of acoustic scenarios remains limited. To more accurately train and evaluate speech separation models for practical use, a dataset that more closely resembles real-world environments is necessary. Specifically, this dataset should include different noise types, cover a wide range of realistic acoustic environments, and have randomly distributed speech overlap ratios.
To address the two issues mentioned above, our main contributions are as follows:
@@
Experiments show that models trained on EchoSet generalized better on real-world data than those trained on benchmark dataset LRS2-2Mix (Li et al., 2023) and Libri2Mix (Cosentino et al., 2020), validating that audio in EchoSet is closer to the physical world. We then comprehensively evaluated TIGER on Libri2Mix, LRS2-2Mix and EchoSet. As the dataset becomes more complex, TIGER’s superiority in performance becomes more significant. On EchoSet, which is the most complicated among the three datasets, TIGER improved the performance by about 5% compared with TF-GridNet, while reducing the parameters and MACs by 94.3% and 95.3% respectively. When tested on real-world data, TIGER also achieved the best separation performance. The remarkable result shows that TIGER provides a new solution for the design of lightweight speech separation models for practical use in the time-frequency domain.
@@ 2 Related Work
Speech separation. Speech separation methods can be divided into time domain and time-frequency domain. Time domain methods directly process the original audio signal. Conv-TasNet (Luo & Mesgarani, 2019) extracts features by temporal convolutional network (Lea et al., 2016). To improve the performance on long sequence data, DPRNN (Luo et al., 2020) divides the temporal sequence into small blocks and alternately performs intra-block and inter-block modeling, which becomes a common paradigm for many following works (Wang et al., 2023; Subakan et al., 2021). Time-frequency domain methods apply a Short-Time Fourier Transform (STFT) to transform the waveform into a joint representation of time and frequency. TF-GridNet (Wang et al., 2023) enhances the temporal context information by a full-band self-attention module. Although TF-GridNet achieved SOTA performance, it entails huge computational costs.
Lightweight models. Some models (Wang et al., 2023; Yang et al., 2022; Subakan et al., 2021) with high computational complexity are difficult to be applied to real-time speech processing on edge devices. To reduce computational costs, TDANet (Li et al., 2023) draws on the attention mechanism of human brains and designs a lightweight structure. In music separation, BSRNN (Luo & Yu, 2023) uses prior knowledge to split band, performing band merging on less important bands to compress the feature while retaining key band information.
Datasets for speech separation. WSJ0-2mix (Hershey et al., 2016) is an early and commonly used fully-overlapping clean speech separation dataset. WHAM! (Wichern et al., 2019) added environmental noises to WSJ0-2mix, and furthermore WHAMR! (Maciejewski et al., 2020) added simple reverberation. Libri2Mix (Cosentino et al., 2020) was proposed based on the observation (Kadıoğlu et al., 2020) that the test performance of Conv-TasNet trained on WSJ0-2mix dropped sharply on other separation datasets. The utterances in Libri2Mix were mixed with sparse overlap, and noises were added to the mixed audio, but reverberation was not considered in Libri2Mix. LRS2-2Mix (Li et al., 2023) was mixed by video clips acquired through BBC. The audio was recorded in real acoustic scenarios, thus containing much noise and reverberation. However, due to the different recording environments of the clips, such as the shapes and materials of the room and objects, the reverberation obtained when the clips were directly mixed is still unrealistic.
@@ 3 TIGER
Figure 1: The overall pipeline of TIGER. We focus on scenarios with only two speakers.
@@ 3.1 Overall Pipeline
The TIGER system (Figure 1) can be divided into five main components: the encoder, the band-split module, the separator, the band-restoration module, and the decoder. Specifically, we first use STFT as the encoder to convert the mixed audio signal $\boldsymbol{S}\in\mathbb{R}^{1\times L}$ into its time-frequency representation $\boldsymbol{X}\in\mathbb{C}^{F\times T}$, where $F$ and $T$ represent the number of frequency bins and time frames, respectively. Next, we apply a frequency band-split strategy, dividing the whole band into $K$ sub-bands of varying widths based on their importance. Each sub-band is transformed into a uniform channel size $N$ using 1D convolutions, and these are then stacked along the frequency dimension to produce the feature representation $\boldsymbol{Z}\in\mathbb{R}^{N\times K\times T}$. Thirdly, $\boldsymbol{Z}$ serves as the input to the separator, which uses FFI blocks with shared parameters to model the acoustic characteristics of each speaker. Subsequently, the band-restoration module restores the sub-bands to the full frequency range using separator output $\boldsymbol{J}_{B}\in\mathbb{R}^{N\times K\times T}$ ($B$ denotes number of blocks in separator), and the mask for each speaker $\boldsymbol{M}_{i}\in\mathbb{C}^{F\times T}$ is applied element-wise product to $\boldsymbol{X}$, producing the separated representation for each speaker $\boldsymbol{H}_{i}\in\mathbb{C}^{F\times T}$. Finally, the inverse STFT is used to generate the clean speech signal $\bar{\boldsymbol{P}}_{i}\in\mathbb{R}^{1\times L}$ for each speaker.
@@ 3.3 Separator
In the separator, the input feature $\boldsymbol{Z}$ passes sequentially through $B$ frequency-frame interleaved (FFI) blocks with shared parameters, as shown in Figure 2. In each FFI block, the frequency path is first used to extract contextual information between different sub-bands, producing $\boldsymbol{Z}_{b,f}\in\mathbb{R}^{N\times K\times T}$. Next, we feed $\boldsymbol{Z}_{b,f}$ into the frame path to further model the contextual information between different time frames, generating $\boldsymbol{Z}_{b,t}\in\mathbb{R}^{N\times K\times T}$.
@@
3.3.1 Multi-scale selective attention module
The MSA module enhances important features through a selective attention mechanism and is divided into three stages: encoding, fusing, and decoding, as shown in Figure 3(a). Taking the MSA module in the frequency path as an example, the input to the module is $\boldsymbol{Z}_{b}$.
The encoding stage. This stage aims to capture multi-scale acoustic features. Specifically, we first use multiple 1D convolutional layers (with a stride of 2 and channel of $H$) to progressively downsample the frequency dimension to $\frac{K}{2^{D}}$, resulting in a set of multi-scale acoustic features $\{\boldsymbol{E}_{d}\in\mathbb{R}^{H\times\frac{K}{2^{d}}\times T}|d=[0,D]\}$, where $d$ denotes the $d$-th layer of downsampling. Next, we apply average pooling layers, denoted as $\lambda(\cdot)$, to downsample all $\boldsymbol{E}_{d}$ to the same frequency resolution $\frac{K}{2^{D}}$. Subsequently, the features with different frequency resolutions are fused into global features $\boldsymbol{G}=\sum_{d=0}^{D}\lambda(\boldsymbol{E}_{d}),\boldsymbol{G}\in\mathbb{R}^{H\times\frac{K}{2^{D}}\times T}$ through addition. Finally, a multi-layer convolutional (MLC) network is used to transform $\boldsymbol{G}$ into $\boldsymbol{G}^{\prime}\in\mathbb{R}^{H\times\frac{K}{2^{D}}\times T}$.
@@ 3.3.1 Multi-scale selective attention module
where $x$ and $z$ represent global features, $y$ represents local features, and $\odot$ denotes element-wise multiplication. This function describes the mathematical process of SA mechanism. We first apply sigmoid function to $x$, generating a value between 0 and 1. Then, the value is used to extract effective features from local information by calculating element-wise product of $\sigma(x)$ and $y$. Finally, we add the product to $z$, fusing global information and filtered local information. In this way, $\{\boldsymbol{L}_{d}\in\mathbb{R}^{H\times\frac{K}{2^{d}}\times T}~{}|~{}d=[0,D]\}$ contains both local and global information, which helps the model better extract the acoustic features in the audio mixture.
The decoding stage. In the $d$-th layer, where $d\in[0,D-1]$, the input consists of the decoding result from the previous layer $d+1$ (denoted as $\boldsymbol{D}_{d+1}\in\mathbb{R}^{H\times\frac{K}{2^{d+1}}\times T}$) and the output $\boldsymbol{L}_{d}$ from the fusing stage at the $d$-th layer. $\boldsymbol{D}_{d+1}$ is processed through the SA module to produce $\boldsymbol{D}_{d}$. Specifically, $\boldsymbol{D}_{d+1}$ is transformed using two 1D convolutions to obtain $\alpha\in\mathbb{R}^{H\times\frac{K}{2^{d+1}}\times T}$ and $\beta\in\mathbb{R}^{H\times\frac{K}{2^{d+1}}\times T}$, while $\boldsymbol{L}_{d}$ is transformed through a 1D convolution to produce $\gamma\in\mathbb{R}^{H\times\frac{K}{2^{d}}\times T}$. We then compute:
@@
where $f$ is defined in equation 5. This formulation integrates the decoding result with the output from the fusing stage to generate the next layer of decoded features. In particular, for the layer where $d=D$, $\boldsymbol{D}_{D}=\boldsymbol{L}_{D}\in\mathbb{R}^{H\times\frac{K}{2^{D}}\times T}$. For the layer where $d=0$, we use one 1D convolution to restore the hidden dimension $H$ in $\boldsymbol{D}_{0}\in\mathbb{R}^{H\times K\times T}$ to the feature dimension $N$, obtaining $\bar{\boldsymbol{Z}}_{b}\in\mathbb{R}^{N\times K\times T}$ as the output of the MSA module. In the MSA module of the frequency path, the frequency dimension $K$ is considered the processing dimension. In the frame path, the time dimension $T$ is considered the processing dimension.